# Bias Behind the Wheel: Fairness Analysis of Autonomous Driving Systems

### v3

@@ Abstract.
This paper analyzes fairness in automated pedestrian detection, a crucial but under-explored issue in autonomous driving systems. We evaluate eight state-of-the-art deep learning-based pedestrian detectors across demographic groups on large-scale real-world datasets. To enable thorough fairness testing, we provide extensive annotations for the datasets, resulting in 8,311 images with 16,070 gender labels, 20,115 age labels, and 3,513 skin tone labels. Our findings reveal significant fairness issues, particularly related to age. The undetected proportions for children are 20.14% higher compared to adults. Furthermore, we explore how various driving scenarios affect the fairness of pedestrian detectors. We find that pedestrian detectors demonstrate significant gender biases during night time, potentially exacerbating the prevalent societal issue of female safety concerns during nighttime out. Moreover, we observe that pedestrian detectors can demonstrate both enhanced fairness and superior performance under specific driving conditions, which challenges the fairness-performance trade-off theory widely acknowledged in the fairness literature. We publicly release the code, data, and results to support future research on fairness in autonomous driving.
Fairness, pedestrian detection, autonomous driving
@@ 1. Introduction
Although significant testing efforts have been made, to the best of our knowledge, the study of fairness testing for autonomous driving systems remains under investigation in the literature. From a Software Engineering (SE) perspective, fairness is considered a non-functional software property, making it an important subject for testing (Zhang et al., 2022). Fairness testing, as an emerging domain within software testing, seeks to uncover fairness issues in software systems (Chen et al., 2024).
@@
To fill the gap, we investigate fairness of eight state-of-the-art Deep Learning (DL)-based pedestrian detectors that have been extensively studied in the research community. Our main focus is to quantitatively assess performance disparities in these detectors across diverse demographic groups, which are widely recognized as group fairness issues (Chen et al., 2022). To enable this investigation, we manually enrich four widely-adopted real-world datasets with gender, age, and skin tone labels, resulting in a collection of 8,311 real-world images annotated with 16,070 gender labels, 20,115 age labels, and 3,513 skin tone labels. Using these labeled datasets, we assess the group fairness of existing pedestrian detectors and also explore how commonly-studied driving scenarios (including various brightness, contrast, and weather conditions) impact the fairness of these detectors.
@@
  • We conduct the first comprehensive study of fairness in autonomous driving systems across various datasets and demographic attributes, evaluating eight widely-studied DL-based pedestrian detectors and uncovering significant fairness issues.
@@ 2.1. Software Fairness
Fairness has gained significant attention in the SE community since its initial exploration by SE researchers in 2008 (Finkelstein et al., 2008). There have been various definitions of fairness in the literature. In this paper, we focus on group fairness, a concept extensively studied in software fairness research (Biswas and Rajan, 2021; Zhang and Harman, 2021; Hort et al., 2021; Chakraborty et al., 2021, 2020; Gohar et al., 2023; Biswas and Rajan, 2020a; Zhang et al., 2023; Zhang and Sun, 2022; Nguyen et al., 2023). Notably, group fairness closely aligns with legal regulations on fairness (Barocas and Selbst, 2016), such as the adherence to the four-fifths rule, a cornerstone of US anti-discrimination law (Biddle, 2005; Feldman et al., 2015; Dwork et al., 2012; Zafar et al., 2017; Wick et al., 2019). Consequently, testing and prioritizing group fairness when building software has emerged as an essential ethical duty and requirement for software engineers (Chakraborty et al., 2021).
@@
In the context of group fairness, certain personal characteristics that require protection against unfairness during decision-making are called sensitive attributes, also known as protected attributes (Zhang et al., 2022; Corbett-Davies and Goel, 2018). Well-recognized sensitive attributes include race, sex, age, pregnancy, familial status, disability status, and more (Chen et al., 2024; Wan et al., 2023). These sensitive attributes typically partition a population into distinct groups: a privileged group and an unprivileged group (Chen et al., 2024). Group fairness entails the equal treatment of these groups by the same machine learning model. However, in practice, members of unprivileged groups often experience systematic disadvantages, resulting from unfair machine learning models. For instance, in the context of a pedestrian detection task, if age is deemed a sensitive attribute, the predictive model may exhibit a bias favoring adult groups over child groups. In this scenario, the adult group is considered the privileged group, while the child group becomes the unprivileged one.
@@
Recently, Chen et al. (Chen et al., 2024) has presented a comprehensive survey of fairness testing research and analyzed its trend. This survey points out that the majority of existing work revolves around tabular data (Biswas and Rajan, 2020b, 2021; Gohar et al., 2023; Zhang and Harman, 2021; Chen et al., 2023). For example, Biswas and Rajan (Biswas and Rajan, 2020b) evaluated fairness of machine learning models on a crowd sourced platform using tabular datasets covering tasks such as credit risk prediction, income prediction, marketing, and loan application. Similarly, Chen et al. (Chen et al., 2023) conducted an empirical study on the group fairness achieved by state-of-the-art bias mitigation methods across eight commonly used tabular data-driven decision tasks. In the contrast, our paper centers on fairness testing for pedestrian detection in autonomous driving systems. We specifically examine three sensitive attributes (i.e., gender, age, and skin tone) that are recognizable in autonomous driving datasets. These sensitive attributes have been demonstrated to be the most widely considered ones in the fairness testing literature (Chen et al., 2024).
@@ 2.2. Autonomous Driving Testing
Autonomous driving is a hot SE research topic, and researchers have proposed various testing techniques for autonomous driving systems (Pei et al., 2017; Tian et al., 2018; Zhang et al., 2018; Zhou et al., 2020; Guo et al., 2022; Zuo et al., 2021). For instance, Tian et al. (Tian et al., 2018) proposed DeepTest, a novel technique using image transformations to emulate potential camera disturbances encountered in driving environments. Zhang et al.  (Zhang et al., 2018) presented DeepRoad, employing Generative Adversarial Networks to craft test images derived from actual driving scenarios. Zhou et al.  (Zhou et al., 2020) introduced DeepBillboard, aiming to generate adversarial billboards that could lead to steering mistakes in autonomous vehicles. Guo et al. (Guo et al., 2022) developed LiRTest, marking the inaugural automated testing technique for LiDAR-equipped autonomous driving software.
While a substantial body of knowledge focuses on assessing the robustness and correctness properties of autonomous driving systems (Zhang et al., 2022), to the best of our knowledge, only a few studies have explored the fairness properties, particularly in the pedestrian detection domain within autonomous driving, indicating that this area remains underdeveloped.
Pedestrian detection is a crucial process that identifies pedestrians within street-level images by providing their predicted locations along with corresponding bounding boxes and confidence scores  (Dollár et al., 2012; Zhang et al., 2016). Despite its significance, research on the fairness of pedestrian detection within autonomous driving is limited. Brandao (Brandao, 2019) explored fairness in pedestrian detection, concentrating only on classic machine learning-based methods. These classic techniques, reliant on manually-defined features, have been eclipsed by deep learning-based pedestrian detection approaches, now prevalent in the autonomous driving industry. Wilson et al.  (Wilson et al., 2019) focused on skin tone bias, confined their analysis to a single dataset and two general object detectors. Similarly, Kogure et al.  (Kogure et al., 2022) explored only age bias using a small-scale dataset and a detection method that is no longer state-of-the-art.
@@
To address this knowledge gap, our paper presents a comprehensive empirical study on revealing fairness issues in pedestrian detection. We conduct experiments using eight popular detection methods and four diverse testing datasets, encompassing different scenarios determined by a variety of factors such as brightness, contrast, and weather conditions. The scale and diversity of our experiments enable us to provide comprehensive insights into the fairness circumstances of the existing pedestrian detectors.
@@ 3.1. Research Questions
Figure 1 illustrates our experimental settings to answer these RQs. In the following, we introduce the pedestrian detectors, benchmark datasets, evaluation metrics, and experimental details.
@@ 3.2. Pedestrian Detectors
In recent years, DL has revolutionized pedestrian detection approaches. We focus our analysis on eight DL-based pedestrian detectors that are widely studied in the autonomous driving community (Zou et al., 2023; Cao et al., 2020). These detectors are pre-trained DL models that researchers and practitioners can directly use for pedestrian detection tasks. They can be classified into two categories (Hasan et al., 2020): general object detectors and pedestrian-specific detectors. Next, we briefly introduce each category and the corresponding pedestrian detectors.
@@
General object detectors YOLOX - MMDetection (MMDetection Developers, nd)
@@
Faster R-CNN X-101-64x4d-FPN
Cascade R-CNN X-101-64x4d-FPN
Pedestrian-specific detectors ALFNet ResNet50 ALFNet (liuwei16, 2018)
CSP ResNet50 Pedestron (hasanirtiza, 2021)
@@
PRNet ResNet50 PRNet (sxlpris, 2020)
General object detectors: General detectors can detect various objects such as cars, traffic lights, and pedestrians. They have great generalization ability but lack pedestrian-specific adaptation (Hasan et al., 2020). They can be categorized into two genres (Zou et al., 2023): two-stage and one-stage detectors. Two-stage detectors propose regions before feature extraction and classification, achieving high accuracy but slower speed; one-stage detectors complete all operations in one step, providing faster speed but lower accuracy. Both types involve trade-offs and are widely used for pedestrian detection. Hence, this paper selects detectors from both categories. For one-stage detectors, we adopt the widely-studied YOLOX (Ge et al., 2021) (a faster extension of the YOLO series (Redmon et al., 2016) used in Apollo’s autonomous driving systems (Apo, 2023)) and RetinaNet (Lin et al., 2017) (which addresses the class imbalance problem). For two-stage detectors, we employ the Faster R-CNN (Ren et al., 2017) (one of the pioneering detectors in the R-CNN family) and Cascade R-CNN (Cai and Vasconcelos, 2018) (which achieves higher accuracy through a cascade of multiple CNNs to refine region proposals). These detectors have been extensively used in the autonomous driving literature (Hasan et al., 2020; Cao et al., 2020; Yu et al., 2020; Braun et al., 2019).
@@ 3.3.1. Dataset Selection
BDD100k gender, age, skin 2,233 day, night
@@ 3.3.2. Sensitive Attribute Labeling
We focus on three sensitive attribute: gender, age, and skin tone. These attributes are identifiable in autonomous driving images and are recognized as the three most extensively studied sensitive attributes in fairness testing literature (Chen et al., 2024). To enable fairness analysis, we need datasets with labels that indicate these sensitive attributes of the humans depicted in the images. Among the datasets investigated herein, the only sensitive attribute already labeled is the skin tone (i.e., light-skin tone and dark-skin tone) for the BDD100k dataset.
@@
The labeling process involves two annotators to minimize the influence of labeling bias. We focus on images that align with the widely-adopted “reasonable subset” principle (Dollár et al., 2012), meaning that we label images containing labeled pedestrians with a height of at least 50 pixels and little to none occlusion. For such images, human annotators can label the sensitive attributes with high confidence (Brandao, 2019). Using the filtered datasets, the two annotators independently label the gender and age attributes for each image. For gender, we follow previous studies (Brandao, 2019) and consider only two labels: male and female. As for age, in line with the literature (Brandao, 2019; Kogure et al., 2022), we classify pedestrians into two labels: child and adult, based on their physical characteristics depicted in the images.
To ensure the reliability of the labeling procedure, both annotators independently label the age and gender attributes for each image. We use Cohen’s Kappa ($\kappa$) (Cohen, 1960), a widely-adopted metric for measuring inter-rater agreement, during the independent labeling process. The obtained $\kappa$ values for age and gender attributes in each of the four datasets are summarized in Table 3. According to the literature (Landis and Koch, 1977), a $\kappa$ value between 0.81 and 1 signifies almost perfect agreement, while a value between 0.61 and 0.8 indicates substantial agreement. In our labeling process, we achieve substantial agreement in gender labeling for the EuroCity-Day dataset and almost perfect agreement for all other tasks. This high level of agreement underscores the reliability of our labeling procedure (Wen et al., 2021; Chen et al., 2021). In cases where the two annotators encounter conflicts, an arbitrator is involved in the discussion to reach a consensus. After the labeling process, the summary of the number of labeled pedestrian instances for each dataset is presented in Table 4.
Table 3. Cohen’s $\kappa$ values for labeling age and gender.
Dataset $\kappa_{age}$ $\kappa_{gender}$
CityPersons 0.847 0.814
EuroCity-Day 0.925 0.800
EuroCity-Night 0.847 0.870
BDD100k 0.828 0.854

### v4

# Bias Behind the Wheel: Fairness Testing of Autonomous Driving Systems

@@ Abstract.
This paper conducts fairness testing of automated pedestrian detection, a crucial but under-explored issue in autonomous driving systems. We evaluate eight state-of-the-art deep learning-based pedestrian detectors across demographic groups on large-scale real-world datasets. To enable thorough fairness testing, we provide extensive annotations for the datasets, resulting in 8,311 images with 16,070 gender labels, 20,115 age labels, and 3,513 skin tone labels. Our findings reveal significant fairness issues, particularly related to age. The proportion of undetected children is 20.14% higher compared to adults. Furthermore, we explore how various driving scenarios affect the fairness of pedestrian detectors. We find that pedestrian detectors demonstrate significant gender biases during night time, potentially exacerbating the prevalent societal issue of female safety concerns during nighttime out. Moreover, we observe that pedestrian detectors can demonstrate both enhanced fairness and superior performance under specific driving conditions, which challenges the fairness-performance trade-off theory widely acknowledged in the fairness literature. We publicly release the code, data, and results to support future research on fairness in autonomous driving.
Fairness testing, pedestrian detection, autonomous driving
@@ 1. Introduction
Although significant testing efforts have been made, to the best of our knowledge, the study of fairness testing for autonomous driving systems remains under investigation in the literature. From the Software Engineering (SE) perspective, fairness is considered a non-functional software property, making it an important subject for testing (Zhang et al., 2022; Sun et al., 2024). Fairness testing, as an emerging domain within software testing, seeks to uncover fairness issues in software systems (Chen et al., 2024a).
@@
To fill the gap, we conduct fairness testing of eight state-of-the-art Deep Learning (DL)-based pedestrian detectors that have been extensively studied in the research community. Our main focus is to quantitatively assess performance disparities in these detectors across diverse demographic groups, which are widely recognized as group fairness issues (Chen et al., 2022). To enable fairness testing, we manually enrich four widely-adopted real-world datasets with gender, age, and skin tone labels, resulting in a collection of 8,311 real-world images annotated with 16,070 gender labels, 20,115 age labels, and 3,513 skin tone labels. Using these labeled datasets, we assess the group fairness of existing pedestrian detectors and also explore how commonly-studied driving scenarios (including various brightness, contrast, and weather conditions) impact the fairness of these detectors.
@@
  • We conduct the first comprehensive study on fairness testing of autonomous driving systems across various datasets and demographic attributes, evaluating eight widely-studied DL-based pedestrian detectors and uncovering significant fairness issues.
@@ 2.1. Software Fairness
Fairness has gained significant attention in the SE community since its initial exploration by SE researchers in 2008 (Finkelstein et al., 2008). There have been various definitions of fairness in the literature. In this paper, we focus on group fairness, a concept extensively studied in software fairness research (Biswas and Rajan, 2021; Zhang and Harman, 2021; Hort et al., 2021; Chakraborty et al., 2021, 2020; Gohar et al., 2023; Biswas and Rajan, 2020a; Zhang et al., 2023; Zhang and Sun, 2022; Nguyen et al., 2023; Chen et al., 2024b). Notably, group fairness closely aligns with legal regulations on fairness (Barocas and Selbst, 2016), such as the adherence to the four-fifths rule, a cornerstone of US anti-discrimination law (Biddle, 2005; Feldman et al., 2015; Dwork et al., 2012; Zafar et al., 2017; Wick et al., 2019). Consequently, testing and prioritizing group fairness when building software has emerged as an essential ethical duty and requirement for software engineers (Chakraborty et al., 2021).
@@
In the context of group fairness, certain personal characteristics that require protection against unfairness during decision-making are called sensitive attributes, also known as protected attributes (Zhang et al., 2022; Corbett-Davies and Goel, 2018; Hort et al., 2024). Well-recognized sensitive attributes include race, sex, age, pregnancy, familial status, disability status, and more (Chen et al., 2024a; Wan et al., 2023). These sensitive attributes typically partition a population into distinct groups: a privileged group and an unprivileged group (Chen et al., 2024a). Group fairness entails the equal treatment of these groups by the same machine learning model. However, in practice, members of unprivileged groups often experience systematic disadvantages, resulting from unfair machine learning models. For instance, in the context of a pedestrian detection task, if age is deemed a sensitive attribute, the predictive model may exhibit a bias favoring the adult group over the child group. In this scenario, the adult group is considered the privileged group, while the child group becomes the unprivileged one.
@@
Recently, Chen et al. (Chen et al., 2024a) have presented a comprehensive survey of fairness testing research and analyzed its trend. This survey points out that the majority of existing work revolves around tabular data (Biswas and Rajan, 2020b, 2021; Gohar et al., 2023; Zhang and Harman, 2021; Chen et al., 2023). For example, Biswas and Rajan (Biswas and Rajan, 2020b) evaluated fairness of machine learning models on a crowd sourced platform using tabular datasets covering tasks such as credit risk prediction, income prediction, marketing, and loan application. Similarly, Chen et al. (Chen et al., 2023) conducted an empirical study on the group fairness achieved by state-of-the-art bias mitigation methods across eight commonly used tabular data-driven decision tasks. In contrast, our paper centers on fairness testing for pedestrian detection in autonomous driving systems. We specifically examine three sensitive attributes (i.e., gender, age, and skin tone) that are recognizable in autonomous driving datasets. These sensitive attributes have been demonstrated to be the most widely considered ones in the fairness testing literature (Chen et al., 2024a).
@@ 2.2. Autonomous Driving Testing
Autonomous driving testing is a hot SE research topic, and researchers have proposed various testing techniques for autonomous driving systems (Pei et al., 2017; Tian et al., 2018; Zhang et al., 2018; Zhou et al., 2020; Guo et al., 2022; Zuo et al., 2021). For instance, Tian et al. (Tian et al., 2018) proposed DeepTest, a novel technique using image transformations to emulate potential camera disturbances encountered in driving environments. Zhang et al.  (Zhang et al., 2018) presented DeepRoad, employing GANs to craft test images derived from actual driving scenarios. Zhou et al.  (Zhou et al., 2020) introduced DeepBillboard, aiming to generate adversarial billboards that could lead to steering mistakes in autonomous vehicles. Guo et al. (Guo et al., 2022) developed LiRTest, marking the first automated testing technique for LiDAR-equipped autonomous driving software.
While a substantial body of knowledge focuses on assessing the robustness and correctness properties of autonomous driving systems (Zhang et al., 2022), to the best of our knowledge, only a few studies have explored the fairness properties, particularly in the pedestrian detection domain within autonomous driving, indicating that this area remains under-explored.
Pedestrian detection is a crucial process that identifies pedestrians within street-level images by providing their predicted locations along with corresponding bounding boxes and confidence scores  (Dollár et al., 2012; Zhang et al., 2016). Despite its significance, research on the fairness of pedestrian detection within autonomous driving is limited. Brandao (Brandao, 2019) explored fairness in pedestrian detection, concentrating on classic machine learning-based methods. These classic techniques, reliant on manually-defined features, have been eclipsed by deep learning-based pedestrian detection approaches, now prevalent in the autonomous driving domain. Wilson et al.  (Wilson et al., 2019) focused on skin tone bias, confined their analysis to a single dataset and two general object detectors. Similarly, Kogure et al.  (Kogure et al., 2022) explored age bias using a small-scale dataset and a detection method that is no longer state-of-the-art.
@@
To address this knowledge gap, our paper presents a comprehensive empirical study on revealing fairness issues in pedestrian detection. We conduct experiments using eight popular DL-based detection methods and four diverse testing datasets, encompassing different scenarios determined by a variety of factors such as brightness, contrast, and weather conditions. We focus on three widely considered sensitive attributes, i.e., gender, age, and skin tone. The scale and diversity of our experiments enable us to provide comprehensive insights into the fairness circumstances of the existing pedestrian detectors.
@@ 3.1. Research Questions
Figure 1 illustrates our experimental settings to answer these RQs. In the following, we introduce the pedestrian detectors, benchmark datasets, evaluation metric, statistical analysis, and experimental details.
@@ 3.2. Pedestrian Detectors
In recent years, DL has revolutionized pedestrian detection approaches. We focus our analysis on eight DL-based pedestrian detectors that are widely studied in the autonomous driving community (Zou et al., 2023; Cao et al., 2020). These detectors are pre-trained DL models that researchers and practitioners can directly use for pedestrian detection tasks. They can be classified into two categories: general object detectors and pedestrian-specific detectors (Hasan et al., 2020). Next, we briefly introduce each category and the corresponding pedestrian detectors.
@@
General object detectors YOLOX - MMDetection (MMD, nd)
@@
Faster RCNN X-101-64x4d-FPN
Cascade RCNN X-101-64x4d-FPN
Pedestrian-specific detectors ALFNet ResNet50 ALFNet (ALF, 2018)
CSP ResNet50 Pedestron (Ped, 2021)
@@
PRNet ResNet50 PRNet (PRN, 2020)
General object detectors: General detectors can detect various objects such as cars, traffic lights, and pedestrians. They have great generalization ability but lack pedestrian-specific adaptation (Hasan et al., 2020). They can be categorized into two categories: two-stage and one-stage detectors (Zou et al., 2023). Two-stage detectors propose regions before feature extraction and classification, achieving high accuracy but slower speed; one-stage detectors complete all operations in one step, providing faster speed but lower accuracy. Both types involve trade-offs and are widely used for pedestrian detection. Hence, this paper selects detectors from both categories. For one-stage detectors, we adopt the widely-studied YOLOX (Ge et al., 2021) (a faster extension of the YOLO series (Redmon et al., 2016) used in Apollo’s autonomous driving systems (Apo, 2023)) and RetinaNet (Lin et al., 2017) (which addresses the class imbalance problem). For two-stage detectors, we employ the Faster RCNN (Ren et al., 2017) (one of the pioneering detectors in the RCNN family) and Cascade RCNN (Cai and Vasconcelos, 2018) (which achieves higher accuracy through a cascade of multiple CNNs to refine region proposals). These detectors have been extensively used in the autonomous driving literature (Hasan et al., 2020; Cao et al., 2020; Yu et al., 2020; Braun et al., 2019).
@@ 3.3.1. Dataset Selection
BDD100k gender, age, skin tone 2,233 day, night
@@ 3.3.2. Sensitive Attribute Labeling
We focus on three sensitive attributes: gender, age, and skin tone. These attributes are identifiable in autonomous driving images and are recognized as the three most extensively studied sensitive attributes in fairness testing literature (Chen et al., 2024a). To enable fairness analysis, we need datasets with labels that indicate these sensitive attributes of the humans depicted in the images. Among the datasets investigated herein, the only sensitive attribute already labeled is the skin tone (i.e., light-skin tone and dark-skin tone) for the BDD100k dataset.
@@
The labeling process involves two annotators to minimize the influence of labeling bias. We focus on images that align with the widely-adopted “reasonable subset” principle (Dollár et al., 2012), meaning that we label images containing labeled pedestrians with a height of at least 50 pixels and little to no occlusion. For such images, human annotators can label the sensitive attributes with high confidence (Brandao, 2019). Using the filtered datasets, the two annotators independently label the gender and age attributes for each image. For gender, we follow previous studies (Brandao, 2019) and consider only two labels: male and female. As for age, in line with the literature (Brandao, 2019; Kogure et al., 2022), we classify pedestrians into two labels: child and adult, based on their physical characteristics depicted in the images.
To ensure the reliability of the labeling procedure, both annotators independently label the gender and age attributes for each image. We use Cohen’s Kappa ($\kappa$) (Cohen, 1960), a widely-adopted metric for measuring inter-rater agreement, during the independent labeling process (Wen et al., 2021; Chen et al., 2021, 2020; Wang et al., 2023a; Liu et al., 2023; Zhang et al., 2024). The obtained $\kappa$ values for gender and age attributes in each of the four datasets are summarized in Table 3. According to the literature (Landis and Koch, 1977), a $\kappa$ value between 0.81 and 1 signifies almost perfect agreement, as shown in Table 4, while a value between 0.61 and 0.8 indicates substantial agreement. In our labeling process, we achieve substantial agreement in gender labeling for the EuroCity-Day dataset and almost perfect agreement for all other tasks. This high level of agreement underscores the reliability of our labeling procedure (Wen et al., 2021; Chen et al., 2021, 2020; Wang et al., 2023a; Liu et al., 2023; Zhang et al., 2024). In cases where the two annotators encounter conflicts, an arbitrator is involved in the discussion to reach a consensus. After the labeling process, the summary of the number of labeled pedestrian instances for each dataset is presented in Table 5.
Table 3. Cohen’s $\kappa$ values for labeling gender and age.
Dataset $\kappa_{gender}$ $\kappa_{age}$
CityPersons 0.814 0.847
EuroCity-Day 0.800 0.925
EuroCity-Night 0.870 0.847
BDD100k 0.854 0.828
Table 4. Cohen’s $\kappa$ values and corresponding agreement levels of inter-rater agreement.
$\kappa$ values Agreement level
$<$ 0 No agreement
0.01-0.20 Slight agreement
0.21-0.40 Fair agreement
0.41-0.60 Moderate agreement
0.61-0.80 Substantial agreement
0.81-1 Almost perfect agreement